OpenStudy (anonymous):

what is the minimum of f(x)=3x^2-12x-1? and how do i find it?

OpenStudy (anonymous):

To find the value of the minimum of the function you can find the value of x that gives that value. The value of x that give the minimum ( or the maximum, if a < 0) on a ax² + bx +c = 0 function it's: \[x _{v}=\frac{ -b }{ 2a }\] So, in that function, you have b = -12 and a = 3. \[x _{v}=\frac{ -(-12) }{ 2*3 } = \frac{ 12 }{ 6 } = 2\] Now you have to calc f(2). \[f(2)=3*(2)^2-12*(2)-1=3*4-24-1=12-24-1=-13\]

OpenStudy (anonymous):

thank you so much....